Problem #1: (15 points) Find a root of 2x3-6r-10 0 Show the results for three iterations of the method.
Answer #1

NEWTON'S METHOD:

Formula:

Xn+1= Xn - (F(Xn) / F'(Xn))

find the approximate root of 2X3-6X-10 upto three iteration

Let F(X)=2X3-6X-10

F'(X)=2(3X2 )-6

=6X2-6

F(0) =2(0)-6(0)-10

=-10

F(1)=2(1)-6(1)-10

= -14

F(2)=2(23)-6(2)-10

= -6

X0 = -4 [ F2 is closer to -4 because 2 to -6 I am taking middle value]

First iteration:

X1= -4 -(-6/90) (substitute x0 value in above formula)

= -4+9.0667

= 5.0667

Second iteration:

X2= 5.0667 - ( (2(5.0667)3-6(5.0667) -10) / (6(5.0667)2-6))

= 5.0667 - ((20.2668-30.4002-10) /(35.4669-6))

= 5.0667 - (20.1334 / 35.4669)

= 4.499

Third iteration:

X3 =4.499 - ((2(4.499)3-6(4.499)-10) / (6(4.499)2-6))

=4.499 - ((182.129-26.994-10) / (121.446-6))

= 4.499 -(145.135 /115.446)

= 4.499 - 1.258

= 3.241